(no subject)

arenton@uvahef.phys.virginia.edu
Tue, 1 Jun 1999 08:00:38 -0400

Shift Summary for owl shift, 01/june/1999

On Shift: Leo B. and Mike A. with Bob T., Elizabeth,
Hogan and Ed present at various times and several
collaborators on the phone.

Hours of beam delivered at all intensities: 5
Hours of beam delivered at nominal intensity: 5
Hours of data (special runs not included):
Hours of special runs (laser scans, muon runs...):3

========================================================================
Runs ended this shift:
13256 and 13257 long mu run with thresholds at 50mV

========================================================================
Accesses during shift:
1 to change DC thresholds back to 40 mV.
1 to power cycle stream 5 repeater in laser rack., this is
same problem seen 24 hours ago.

On both these accesses there were problems with rad monitor
trips that had to be reset.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

For a shift that originally was supposed to have no beam, it
was very eventful, both good and bad.

The good was at the beginning. DC plateau's at 50mV threshold
were finished and a long run was taken at the final settings.
This was runs 12356 which ended when a disk filled up, and 12357.
At 2:00 Hogan and Elizabeth made an access to set DC thresholds
back to 40 mV and to make minor changes in photon veto HV's
for gain balancing study.

After this NM4AN was turned on at 1600A. A new run hung with
the repeater problem seen earlier, an access was required to
power cycle the repeater in the laser rack.

at 2:10 ONSPLIT asked for mount of tape QKMU01, which was done.

Then we tried to take e832_ecal data. The CsI data looked fine,
the beam was well centered, etc. There were serious problems
with the tracking, however. There may be two problems. First
the DC raw time distributions looked very ratty with hadron
beam. (However the time distributions for hits on tracks were
OK). The SOD distributions were somewhat off nominal.
The second problem was that although tracks and vertices were
being found they all appeared (in the display) to have infinite
momentum, and no tracks got into E/p or mass plots. Naturally
this leads one to suspect that the pt-kick was was wrong, but
after several hours of searching it is not yet clear if and why
this is so.

There were a number of CsI laser RMS too wide warnings, which
went away on their own. This appeared worse during hadron beam.

The 3x Kumquats hung several times, delaying run starts.

Beam went away at 5:00, due to safety system problems in
the Main Injector, and didn't return by the end of the shift.